### v2

# High-Order Hermite Optimization: Fast and Exact Gradient Computation in Open-Loop Quantum Optimal Control using a Discrete Adjoint Approach

Abstract
This work introduces the High-Order Hermite Optimization (HOHO) method, an open-loop discrete adjoint method for quantum optimal control. Our method is the first of its kind to efficiently compute exact (discrete) gradients when using continuous, parameterized control pulses while solving the forward equations (e.g. Schrödinger’s equation or the Linblad master equation) with an arbitrarily high-order Hermite Runge-Kutta method. The HOHO method is implemented in QuantumGateDesign.jl an open-source software package for the Julia programming language, which we use to perform numerical experiments comparing the method to Juqbox.jl. For realistic model problems we observe speedups up to 775x.
journal: Journal of Computational Physics
[…]
[label1]organization=Michigan State University, Department of Computational Mathematics, Science, and Engineering & Department of Physics and Astronomy,
[…]
[label2]organization=Department of Mathematics, Virginia Tech, addressline=, city=Blacksburg, postcode=24061, state=VA, country=USA
1 Introduction
Quantum optimal control (QOC) is the theory and practice of manipulating quantum mechanical systems. Quantum technologies, including quantum computing, quantum information processing, quantum simulation, and quantum sensing, depend on the ability to precisely manipulate quantum systems. We focus on the application of QOC to quantum computing algorithms. Such quantum computing algorithms are implemented by applying a series of quantum gates consisting of unitary transformations of a quantum system over a fixed time duration.
A quantum gate is realized in hardware by varying the amplitude of control pulses in time: microwave pulses in the case of superconducting circuits [21, 33], and laser pulses in the case of trapped ions [35] or neutral atoms [24, 37]. These control pulses may be generated experimentally using arbitrary waveform generators (AWGs) for microwaves, or acoustic/electro-optic modulators for lasers. In this work we will focus on superconducting circuits, although the methods we describe may be applied to a variety of quantum systems.
For certain quantum gates operating on small, simple systems, analytic control pulses can be found. For larger, more practically useful systems, numerical optimization is used to find control pulses which implement a desired quantum gate. Gate fidelity is measured by an objective function and the control pulses are parameterized by a set of control parameters.
QOC methods may be open-loop, in which case the quantum system is simulated numerically, or closed-loop, in which case feedback from real experiments is used. For the problem of gate design, closed-loop typically means applying control pulses and measuring the performance of the gate at the end of the pulse application (e.g. by randomized benchmarking [27] or quantum state tomography [20]). In general, open-loop methods excel at accurately computing gradients of the objective function with a cost that does not scale with the number of control parameters, whereas closed-loop methods excel at correctly characterizing the behavior of the the gate by using experimental results instead of simulated results based on potentially inaccurate theoretical models, and at scaling with the number of qubits. Control pulses found using open-loop control may be further tuned using closed-loop control, so even when closed-loop control is necessary for the final implementation of a gate, open-loop control is still a valuable tool.
The adjoint method is a technique for efficiently calculating the gradient of an objective functional that depends on the solution of a differential equation [13]. By “efficiently”, we mean in particular that the computational cost comes mainly from solving two initial value problems (the forward and adjoint equations), and does not depend significantly on the number of control parameters. Within this framework, there are continuous adjoint methods, in which the adjoint equations are derived from the continuous forward forward equations (optimize then discretize), and discrete adjoint methods, in which the adjoint equations are derived from the discretization of the forward equations by a numerical method (discretize then optimize), so that the gradients computed are exact with respect to the numerical method used. This is especially important when optimizing the objective functional using quasi-Newton methods such as L-BFGS [30], because even modest errors in the gradient may ruin the Hessian approximation, so that the advantage of using curvature information is lost.
The primary result of this work is the introduction of the High-Order Hermite Optimization (HOHO) method for QOC, an open-loop discrete adjoint method which is the first of its kind to efficiently compute exact (discrete) gradients when using continuous, parameterized control pulses while solving the forward equations (e.g. Schrödinger’s equation or the Linblad master equation) with an arbitrarily high-order numerical method. Namely, we solve the forward equations using Hermite Runge-Kutta methods, and use their relatively simple form when applied to systems of ODEs to derive a procedure that works for every order version of the method. Additionally, we present techniques that we use to make the method more computationally efficient for the specific models and parameterizations found in QOC. The HOHO method is implemented in QuantumGateDesign.jl [22], an open-source software package for the Julia programming language [4].
The most ubiquitous methods for open-loop QOC are Gradient Ascent Pulse Engineering (GRAPE) [11] and Krotov’s method [12]. These methods rely on the assumption that the control pulses are piecewise-constant, which allows the forward equations for be solved using matrix exponentiation for each time interval over which the control pulses are constant. Consequently, the discretized forward equations are analytically simple, which makes it easy to derive exact expressions for partial derivatives of the objective function. GRAPE is a discrete adjoint method, and updates all the control parameters at once using the gradient, while Krotov’s method updates only a single control parameter at a time, with the advantage that the update scheme is guaranteed to converge monotonically.
The piecewise-constant control pulse assumption has several drawbacks. First, the number of control parameters is exactly the number of timesteps taken, so that longer or more “expressive” gates result in a more challenging optimization problem in addition to a more challenging numerical simulation problem. Second, because each control pulse amplitude corresponds to a different control parameter, the amplitude may vary significantly between time intervals, resulting in “high-bandwidth” control pulses that are difficult to realize experimentally on an AWG. Third, matrix exponentiation scales poorly with system size, compared to other numerical timestepping methods.
The first and second drawbacks of the piecewise-constant control pulse assumption have essentially been solved by the Gradient Ascent in Function Space (GRAFS) method [25], which still assumes the control pulses are piecewise-constant, but parameterizes them in terms of a set of basis functions to enable high-resolution control pulses while drastically reducing the number of control parameters. GRAFS uses a modified version of the gradient computation procedure used in GRAPE to keep the gradient computation efficient. To comply with experimental constraints, Slepian sequences [36] are used to synthesize bandwidth-limited controls.
Before GRAFS, The idea of parameterizing the control pulses in terms of basis functions was also explored in the Chopped Random Basis (CRAB) method [6], which uses smooth, parameterized control pulses. But as a tradeoff lacks a way to compute the gradient (except by finite-difference approximation), and instead uses a direct search to optimize the objective function. The Gradient Optimization of Analytic Controls (GOAT) method [26] also uses smooth, parameterized control pulses, but performs a gradient-based search by differentiating the forward equation with respect to each control parameter (forward-mode differentiation). The GOAT method produces exact gradients, but must solve an initial value problem for every control parameter, so that the gradient computation is not efficient. We also mention the Padé Integrator COllocation method (PICO) [38] which optimizes over the quantum states directly (a direct method), implementing the Schrödinger’s equation as a constraint in the optimization instead of having it enforced implicitly by a numerical timestepping scheme.
Recently, advances have been made in using discrete adjoint methods with smooth, parameterized control pulses to efficiently compute the gradient when using second-order methods to solve the forward equations. Respectively, Juqbox.jl [2] does QOC for closed quantum systems using Störmer-Verlet as the timestepping method, and Quandary [18] does QOC for open quantum systems using the implicit midpoint method as the timestepping method.
In principle, efficient gradient computation can be achieved for any method of solving the forward equations by using reverse-mode automatic differentiation (AD) [16, 28]. Indeed, AD has been applied in the field of of QOC [23, 1, 34], and works similarly to the discrete adjoint method. The main difference between the two methods is that the discrete adjoint method results in clear, analytic forward and adjoint equations to be solved, which allows for the development of more efficient algorithms and software than is typically implemented by a pure AD approach.
In particular, reverse-mode AD must store the expression graph and the values of intermediate variables in that graph. This results in memory storage that scales the system size and with the number of matrix-vector multiplications (which scales with the number of timesteps taken, and hence also with the stiffness of the system) and makes optimal control of large quantum systems, especially open quantum systems, infeasible. In contrast, discrete adjoint methods are often able to store intermediate in a small number of pre-allocated arrays, drastically reducing the total memory required. This also drastically reduces the number of large (the size of the system) dynamic memory allocations required, which can have a significant impact on performance when solving sparse differential equations (which typically appear in QOC), since the costs of the memory allocation and iteratively solving a linear system of equations (e.g. in an implicit method) at each timestep may both be $\mathcal{O}(N)$, where $N$ is the size of the system. Additionally, the discrete adjoint method differentiates each discretized forward equation to produce a corresponding adjoint equation. This allows each both the forward and adjoint equations to be solved iteratively. In contrast, in most software implementations of AD will differentiate through each step of the iterative solver, resulting in increased memory usage and unnecessary computations compared to the discrete adjoint method.
Some of the shortcomings of AD are addressed by the semi-automatic differentiation approach [14]. In this approach, the objective functional is rewritten in terms of intermediary variables. By carefully choosing the intermediate variables the computational graph for can be kept small, reducing the total memory required compared to automatic differentiation, while still keeping the flexibility in the choice of the objective functional.
With the HOHO method, we aim to improve the state-of-the-art in discrete adjoint QOC methods using smooth, parameterized control pulses by using arbitrarily high-order numerical methods, enabling the optimal control of stiff quantum systems, which would be extremely computationally expensive to simulate using lower-order methods.
The rest of the paper is organized as follows. In Section 2, we provide a mathematical description of quantum computing and quantum optimal control. In Section 3, we describe the Hermite Runge-Kutta methods we use in our numerical simulations, and derive a discrete adjoint method for efficiently computing the gradient of arbitrary objective functionals for arbitrarily high-order Hermite Runge-Kutta methods. This is the main novel result of this work. In Section 4.1, we perform numerical experiments on a Rabi oscillator problem to verify the correctness of our method and its implementation. In Section 4.2, we use our method to solve a gate design problem for a tripartite system, showcasing the speed of our method. Finally, in Section 5, we provide concluding remarks and propose ideas for future work.
2 Problem Description
We now present a more precise description of quantum computing and QOC before detailing the HOHO method.
2.1 Quantum Computing
The state of a closed, quantum system is represented by a complex-valued state vector [^3] In quantum computing literature, states are often represented using bra ket notation, in which case the state $\boldsymbol{\psi}\in\mathbb{C}^{N}$ corresponds to $|\psi\rangle$ in a particular basis. Note that $|\psi\rangle$ only represents a physical state: it does not choose a basis to use to represent that state. We switch between the two notations when convenient, but mostly stick to representing states as complex-valued vectors. $\boldsymbol{\psi}(t)\in\mathbb{C}^{N}$. Each observable (a quantity which can be physically measured) of the system is associated with a linear operator. The eigenvalues of the operator are the possible outcomes of the measurement, and the corresponding eigenvectors are the states for which that measurement will be observed.
Each entry of the state vector corresponds to an observable state, and the magnitude of the entry squared gives the probability to observe the corresponding state upon measurement. Consequently, the length of the state vector is $\|\boldsymbol{\psi}(t)\|_{2}=1$ at all times because the probabilities of observing each possible outcome must sum to one. This also implies that the time evolution of the state vector is unitary. A state which has a nonzero magnitude for two or more states is said to be in a superposition of those states.
In this work we consider only closed quantum systems. In a quantum computing context, this means considering only time scales short enough that there is no significant interaction between the quantum computer and the environment.
2.2 Governing Equation: Schrödinger’s Equation
The time evolution of the state vector from an initial state $\boldsymbol{\psi}_{0}$ in a closed quantum system is governed by Schrödinger’s equation [^4] We always choose our units so that $\hbar=1$.
$\frac{d}{dt}\boldsymbol{\psi}(t)=-iH(t)\boldsymbol{\psi}(t),\quad\boldsymbol{\psi}(0)=\boldsymbol{\psi}_{0}\in\mathbb{C}^{N},$ (1)
where $H(t)\in\mathbb{C}^{N\times N}$ is the Hamiltonian of the system. We assume that the Hamiltonian is controlled by a function parametrized by ${N_{\rm P}}$ values collected in the control vector $\boldsymbol{\theta}\in\mathbb{R}^{{N_{\rm P}}}$. Accordingly, we write Schrödinger’s equation as
$\frac{d}{dt}\boldsymbol{\psi}(t)=-iH(t;\boldsymbol{\theta})\boldsymbol{\psi}(t),\quad\boldsymbol{\psi}(0)=\boldsymbol{\psi}_{0}\in\mathbb{C}^{N}.$ (2)
We make an additional assumption that the Hamiltonian can be decomposed into drift and control Hamiltonians
$H(t;\boldsymbol{\theta})=H_{d}+H_{c}(t;\boldsymbol{\theta}),$ (3)
where the drift Hamiltonian governs the behavior of the system if no control is applied.
Quantum computing systems consists of several qubits whose individual states (when not entangled) are represented by a state vector $\boldsymbol{\psi}_{qubit}\in\mathbb{C}^{2}$. In bra-ket notation, the standard basis elements of $\mathbb{C}^{2}$ are written as ${\mathbf{e}}_{1}=|0\rangle$ and ${\mathbf{e}}_{2}=|1\rangle$. The vector $|0\rangle$ is called the ground state of a qubit, and $|1\rangle$ is called the excited state of a qubit (or first excited state, if the qubit has more than two levels, in which case the qubit is called a qudit).
A collection of $Q$ (possibly entangled) qubits is described by a state vector $\boldsymbol{\psi}(t)\in\mathbb{C}^{2^{Q}}$. The basis elements may be constructed using Kronecker products of the state vectors of each qubit. E.g. the state corresponding to a 2 qubit system with the first qubit in the ground state and the second qubit in the excited state is $|01\rangle={\mathbf{e}}_{1}\otimes{\mathbf{e}}_{2}\in\mathbb{C}^{2^{2}}$. The standard ordering of the basis elements in bra-ket notation is $|00\rangle,|01\rangle,|10\rangle,|11\rangle$. A superposition of observable states is represented by a linear combination of the standard basis elements. For example $(|0\rangle+|1\rangle)/\sqrt{2}$ corresponds to a qubit in a superposition of the ground and excited states. Upon measurement, there is a 50/50 chance of observing the system to be in either state.
Hamiltonians are often given in terms of raising/lowering operators. For a single $n$-level qudit, the lowering operator takes the form
$a\coloneqq\begin{bmatrix}0&\sqrt{1}&&\\ &\ddots&\ddots&\\ &&\ddots&\sqrt{n-1}\\ &&&0\\ \end{bmatrix}\in\mathbb{R}^{n\times n},$
, and the raising operator is given by $a^{\dagger}$. For a system of $Q$ qudits, with the $q$-th qudit having $n_{q}$ levels, the lowering operator of the subsystem corresponding to the $q$-th qudit is given by
$a_{q}\coloneqq I_{n_{Q}}\otimes\cdots\otimes I_{n_{q}+1}\otimes a_{q}\otimes I_{n_{q}-1}\otimes\cdots\otimes I_{n_{1}},\quad N=\prod_{q=1}^{Q}n_{q}.$
As a representative example, for a system of $Q$ transmon qubits in the dispersive limit the drift Hamiltonian is
$H_{d}=\sum_{q=1}^{Q}\left(\omega_{q}a_{q}^{\dagger}a_{q}-\frac{\xi_{q}}{2}a_{q}^{\dagger}a_{q}^{\dagger}a_{q}a_{q}-\sum_{p>q}\xi_{pq}a_{p}^{\dagger}a_{p}a_{q}^{\dagger}a_{q}\right).$ (4)
Here $\omega_{q}$ and $\xi_{q}$ are the ground state transition frequency and self-Kerr coefficient, respectively, of the $q$-th qudit, and $\xi_{pq}$ is the cross-Kerr coefficient between the $p$-th and $q$-th qudits. The self-Kerr coefficients are typically (on IBM devices) an order of magnitude smaller than the ground state transition frequencies, and the cross-Kerr coefficients are typically 2-3 orders of magnitude smaller than the ground state transition frequencies.
2.3 Optimal Control Problem
We are interested in optimal control problems where the goal is to control a dynamic system over a time interval $t\in[0,T]$ to minimize some objective function, which may depend on the state of the dynamic system both at the final time $t=T$ and throughout the duration $t\in[0,T]$. The two most common problems in quantum optimal control are state-to-state transfer and gate design problems. In this paper we focus on gate design problems.
Algorithms for quantum computers are formulated in terms of quantum logic gates, which are unitary operators used to perform logical operations on subsystems of the quantum computer which consist of one or more qubit(s). Examples of single qubit gates include the Pauli-X/Y/Z and Hadamard gates [29]. Two qubit gates include the CNOT (controlled NOT) and SWAP gates. An example of a three qubit gate is the Toffoli gate. A gate is a unitary operator, so in any given basis the gate can be defined as a unitary matrix. Gate design is the problem of finding controls (parameterized by $\boldsymbol{\theta}$) which result in a Hamiltonian $H(t;\boldsymbol{\theta})$ that implements the action of a desired gate in time $T$.
More precisely, we control the time evolution of $E$ state vectors [^5] For $n$ qubits, $E=2^{n}$. When each qubit is modeled as a pure two-level system, then $N=E$. If additional levels are modeled, then $N>E$. that form a basis of the computational subspace:
$\frac{d}{dt}U(t)=-iH(t;\boldsymbol{\theta})U(t),\quad U(0)=U_{0}\in\mathbb{C}^{N\times E}.$ (5)
A state-to-state transfer problem is equivalent to a gate design problem where $E=1$.
[…]
$\mathcal{F}(U(T;\boldsymbol{\theta}),U_{\text{target}})\coloneqq\frac{1}{E^{2}}\left|\left\langle U(T;\boldsymbol{\theta}),U_{\text{target}}\right\rangle_{F}\right|^{2}.$ (6)
This definition has several motivations. When $U(T;\boldsymbol{\theta})$ and $U_{\text{target}}$ are both unitary, we have $0\leq\mathcal{F}\leq 1$. Specifically, $\mathcal{F}=1$ when $U(T;\boldsymbol{\theta})$ and $U_{\text{target}}$ are equal, and $\mathcal{F}=0$ when the columns of $U(T;\boldsymbol{\theta})$ and $U_{\text{target}}$ are orthogonal (i.e. $U(T;\boldsymbol{\theta})$ transforms each initial state to a state that is orthogonal to its target state). We also have that $\mathcal{F}$ is convex with respect to $U(T;\boldsymbol{\theta})$ and that the value of $\mathcal{F}$ is invariant to changes in global phase:
$\mathcal{F}(U(T;\boldsymbol{\theta}),U_{\text{target}})=\mathcal{F}(e^{i\delta_{1}}U(T;\boldsymbol{\theta}),e^{i\delta_{2}}U_{\text{target}}),\quad\ \forall\ \delta_{1},\delta_{2}\in\mathbb{R}.$
This last property is physically important, as the global phase of a state is not physically observable, since we take the magnitude squared of the elements of a state vector when calculating the probabilities of measurement outcomes.
It is common for optimization to expect an optimization problem to be formulated as a minimization problem, so instead of maximizing the fidelity, we often consider the equivalent problem of minimizing the trace infidelity
$\mathcal{I}(U(T;\boldsymbol{\theta}),U_{\text{target}})=1-\mathcal{F}(U(T;\boldsymbol{\theta}),U_{\text{target}}).$ (7)
To illustrate the idea of gate design, we provide the example of implementing Hadamard gate for a system described by the Hamiltonian
$H(t)=\frac{\omega_{0}}{2}\sigma_{z}+c(t)\sigma_{x},\quad\textrm{where }\sigma_{z}\coloneqq\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},\quad\sigma_{x}\coloneqq\begin{bmatrix}0&1\\ 1&0\end{bmatrix},\quad\omega_{0}=0.1\textrm{ GHz},$
by shaping the control pulse $c(t)$ to reach a gate infidelity less than 0.01%. The resulting pulse and the evolution of the population of each state starting from each initial condition is shown in Figure 1.
Figure 1: Control pulse implementing the Hadamard gate for the Hamiltonian $H(t)=0.5\omega_{0}\sigma_{z}+c(t)\sigma_{x}$.
The infidelity is not necessarily bounded below by 0 when a numerical method is used to approximate $U(T;\boldsymbol{\theta})$. This is because the numerical approximation $\tilde{U}(T;\boldsymbol{\theta})\approx U(T;\boldsymbol{\theta})$ is not necessarily unitary. As an extreme example, if the inner product $\left\langle U(T;\boldsymbol{\theta}),U_{\text{target}}\right\rangle_{F}$ is nonzero, the infidelity can become arbitrarily negative and large if the numerical approximation is $\tilde{U}(t;\boldsymbol{\theta})=cU(t;\boldsymbol{\theta})$, with $c\gg 1$. Consequently, an optimizer may try to minimize the infidelity (7) by increasing the spectral norm of $\tilde{U}(T;\boldsymbol{\theta})$, which is a completely numerical and non-physical phenomenon. To combat this, we may instead use the generalized infidelity [32]
$\mathcal{I}_{G}(U(T;\boldsymbol{\theta}),U_{\text{target}})\coloneqq\frac{1}{E}\left\lVert U(T;\boldsymbol{\theta})\right\rVert_{F}^{2}-\frac{1}{E^{2}}\left|\left\langle U(T;\boldsymbol{\theta}),U_{\text{target}}\right\rangle_{F}\right|^{2},$ (8)
which equals the standard infidelity when $U(T;\boldsymbol{\theta})$ is unitary, but is bounded between 0 and 1 even for non-unitary $U(T;\boldsymbol{\theta})$. Increasing the spectral norm of $U(T;\boldsymbol{\theta})$ increases $\left\lVert U(T;\boldsymbol{\theta})\right\rVert_{F}^{2}$ (because $\left\lVert A\right\rVert_{2}\leq\left\lVert A\right\rVert_{F}$), which deters an optimizer from exploiting this non-physical phenomenon to minimize the infidelity.
We use either the trace infidelity or the generalized infidelity as the main component of our objective function, but we may add additional terms in order to achieve other desirable properties. In our case, we also add a term measuring the population of the “guard” states (which are not in the computational subspace) throughout the duration of the gate:
[…]
Here $W$ is an $N\times N$ matrix whose null space is the computational subspace. This is done to motivate the truncation of the Hilbert space, which for most quantum computing architectures has infinite dimensions. By including the guard penalty (9) in our objective function, we prevent “leakage” of the state vector into the non-computational subspace. Physically, this means punishing the population of highly excited states. In our models, we typically include one or two guard levels for each qubit. Then, in the computational basis, $W$ is diagonal, with $W_{i,i}=1$ when the $i$-th basis state is an essential state. For a non-essential state $j$, we have $W_{j,j}=w_{j}\neq 0$, where the weight $w_{j}$ can be adjusted to suppress the population of certain guard levels more than others.
2.4 Effect of State Errors on Fidelity
The error, $\mathcal{E}(\delta,\Delta t)$ in the state vector, caused by a perturbation $\delta$ in a coefficient of the control, or the truncation error $\Delta t^{p}$ for a $p$-th order accurate numerical method, will be $\mathcal{E}(\delta,\Delta t)=\mathcal{O}(C\delta+\tilde{C}\Delta t^{p})$. Here $C$ and $\tilde{C}$ are constants that can depend on the solution and the choice of numerical method but not on $\delta$ and $\Delta t$. Here the goal is to determine the control accurately and it is therefore important that $\tilde{C}\Delta t^{p}\ll C\delta+\tilde{C}$, so that the optimization landscape depends more heavily on the real, physical change in the state vector due to the perturbation $\delta$ than on the non-physical change due to the truncation error. Clearly this is easier to achieve for a high order method.
We note, however, that the error in the infidelity is quadratic in the state error. As a result, the state error can generally be allowed to be quadratically larger than the error in the fidelity. For other objectives, the error may be linear in the state error, and hence a higher level of accuracy in the state will be required.
As an example, we optimize of the generalized infidelity of a CNOT gate for a system of two qubits coupled to a resonator bus (which will be described in more detail in Section 4.2). The optimization is performed using a fourth-order method to solve Schrödinger’s equation numerically, with the timestep size chosen relative final state error is approximately $10^{-1}$. For each optimization iteration, we solve the same equation again, choosing the timestep size so that the relative final state error is approximately $10^{-7}$ In Figure 2, we plot the generalized infidelity as computed by the low-accuracy and high-accuracy solutions, as well as the relative final state error in the low-accuracy solution, as a function of the number of optimization iterations. We see that although after 500 iterations the optimization produces a gate with a generalized infidelity of approximately $10^{-4}$ as computed by the low-accuracy method, the actual generalized infidelity is closer to $10^{-1}$, and the optimization iterations performed after the generalized infidelity had reached $10^{-2}$ were wasted computations which actually increased the generalized infidelity.
Figure 2: Generalized infidelity of a CNOT gate, as computed by a low-accuracy and high-accuracy numerical method as a function of the number of optimization iterations, where the optimization is performed using the low-accuracy method.
3 Numerical Methods
In what follows, we convert Schrödinger’s equation (1) into its real-valued formulation, which is commonly done in numerical software. The real-valued form is
$\begin{bmatrix}\frac{d{\mathbf{u}}}{dt}\\ \frac{d{\mathbf{v}}}{dt}\end{bmatrix}=\begin{bmatrix}S(t)&K(t)\\ -K(t)&S(t)\end{bmatrix}\begin{bmatrix}{\mathbf{u}}\\ {\mathbf{v}}\end{bmatrix},$
where we have defined ${\mathbf{u}}\coloneqq{\rm Re}\,(\boldsymbol{\psi})$, ${\mathbf{v}}\coloneqq{\rm Im}\,(\boldsymbol{\psi})$, $K(t)\coloneqq{\rm Re}\,(H(t))$, and $S(t)\coloneqq{\rm Im}\,(H(t))$. Because $H(t)$ is Hermitian, we have $K(t)^{T}=K(t)$ is symmetric, and $S(t)^{T}=-S(t)$ is anti-symmetric. For convenience, we also define
${\mathbf{w}}\coloneqq\begin{bmatrix}{\mathbf{u}}\\ {\mathbf{v}}\end{bmatrix},\quad A(t)\coloneqq\begin{bmatrix}S(t)&K(t)\\ -K(t)&S(t)\end{bmatrix}$
so that we can write the real-valued formulation of Schrödinger’s equation more compactly as
$\frac{d}{dt}{\mathbf{w}}(t)=A(t){\mathbf{w}}(t).$ (10)
3.1 Hermite Runge-Kutta Methods
We now describe Hermite Runge-Kutta methods [7, 17] and how they can be used to efficiently compute discrete adjoints. We follow the description in [7] which explicitly considered the case of a variable coefficient linear system of ordinary differential equations and thus can be directly applied to the problem at hand.
As discussed above, the models we are interested in take the form
$\frac{d}{dt}{\mathbf{w}}(t)=A(t;\boldsymbol{\theta}){\mathbf{w}}(t)\quad 0\leq t\leq T,\quad{\mathbf{w}}(0)={\mathbf{w}}_{0}\in\mathbb{R}^{N}.$ (11)
Here the matrix $A(t;\boldsymbol{\theta})$ depends on time and a vector $\boldsymbol{\theta}$ which parameterizes the the time-dependence of the controls. For each simulation of (11) the parameter vector $\boldsymbol{\theta}$ will be fixed and for notational convenience we suppress it in this sub-section and write $A(t;\boldsymbol{\theta})=A(t)$.
To approximate solutions to (11) we introduce the equidistant grid $t_{n}=n\Delta t,\,n=0,\ldots,{N_{\rm T}},\,\Delta t=T/{N_{\rm T}}$. We denote the numerical approximation to the the solution ${\mathbf{w}}(t_{n})$ by ${\mathbf{w}}_{n}$.
Now, to compute ${\mathbf{w}}_{n+1}$ from initial data ${\mathbf{w}}_{n}$ we first recast (11) as
[…]
replace ${\mathbf{w}}(t_{n})$ with ${\mathbf{w}}_{n}$, and then approximate the integral by quadrature.
One way of constructing a Runge-Kutta method is to introduce $s$ nodes between $t_{n}$ and $t_{n+1}$ and then integrate the interpolant of $\frac{d{\mathbf{w}}(t)}{dt}$ exactly. This is also the approach in a Hermite Runge-Kutta method where the $s=p+q$ nodes are split into two sets of nodes containing $p$ and $q$ nodes respectively. Unlike most Runge-Kutta methods, here there are no nodes in the interior of the interval $[t_{n},t_{n+1}]$ and instead $p$ nodes coalesce at $t_{n}$ and $q$ nodes coalesce at $t_{n+1}$. Then the interpolation polynomial becomes a Hermite interpolation polynomial of degree $p+q-1$. That is
$\frac{d{\mathbf{w}}(t)}{dt}\approx\sum_{j=0}^{p-1}\mathfrak{L}_{0,j}(t)\frac{d^{j+1}{\mathbf{w}}(t_{n})}{dt^{j+1}}+\sum_{j=0}^{q-1}\mathfrak{L}_{1,j}(t)\frac{d^{j+1}{\mathbf{w}}(t_{n+1})}{dt^{j+1}},$
where $\mathfrak{L}_{0,j}(t)$ and $\mathfrak{L}_{1,j}(t)$ are the unique degree $p+q-1$ polynomials (sometimes called generalized Lagrange polynomials, see [9]) defined through the conditions
$\frac{d^{k}\mathfrak{L}_{0,j}(t_{n})}{dt^{k}}=\delta_{kj},\quad\frac{d^{l}\mathfrak{L}_{1,j}(t_{n+1})}{dt^{l}}=\delta_{lj},\quad\frac{d^{l}\mathfrak{L}_{0,j}(t_{n+1})}{dt^{l}}=0,\quad\frac{d^{k}\mathfrak{L}_{1,j}(t_{n})}{dt^{k}}=0,\\ j=0,\dots,p+q-1,\quad k=0,\dots,p-1,\quad l=0,\dots,q-1.$
After evaluating the integrals of $\mathfrak{L}_{0,j}(t)$ and $\mathfrak{L}_{1,j}(t)$ the resulting one-step method [7] takes the form
$\sum_{j=0}^{q}(-1)^{j}c_{j}^{qp}\frac{{\mathbf{w}}_{n+1}^{(j)}}{j!}\Delta t^{j}=\sum_{j=0}^{p}c_{j}^{pq}\frac{{\mathbf{w}}_{n}^{(j)}}{j!}\Delta t^{j},$ (12)
where we have introduced the shorthand notation ${\mathbf{w}}_{n}^{(j)}=\frac{d^{j}{\mathbf{w}}(t_{n})}{dt^{j}}$. Here the coefficients $c_{j}^{pq}$ can be computed explicitly:
$c_{j}^{pq}=\binom{p}{j}\bigg{/}\binom{p+q}{j}=\frac{p!(p+q-j)!}{(p+q)!(p-j)!}.$
For the case of a closed quantum systems we restrict ourselves to methods with $p=q$, for which this method is A-stable. But we note that for open quantum systems with subsystems that decohere fast, L-stable methods with $q>p$ could be explored.
3.2 Computing Higher Order Derivatives and Solving The Linear System of Equations
For a general non-linear ODE the computation of the derivatives ${\mathbf{w}}_{n}^{(j+1)}$ can be cumbersome but when the ODE is governed by (11) this computation can be done efficiently and recursively [7, 19] by differentiating (11) $j$ times and using the generalized Leibniz rule:
${\mathbf{w}}_{n}^{(j+1)}=\sum_{i=0}^{j}\binom{j}{i}A_{n}^{(j-i)}{\mathbf{w}}_{n}^{(i)},\quad j=0,\dots,q-1.$ (13)
If we define the linear operators
$D_{n,0}=I,\quad D_{n,j+1}=\sum_{i=0}^{j}\binom{j}{i}A_{n}^{(j-i)}D_{n,i},\quad j=0,\dots,q-1,$ (14)
we observe that ${\mathbf{w}}_{n}^{(j)}=D_{n,j}{\mathbf{w}}_{n}$, and hence the method (12) can be rewritten as
$\left(\sum_{j=0}^{q}(-1)^{j}c_{j}^{qp}\frac{1}{j!}\Delta t^{j}D_{n+1,j}\right){\mathbf{w}}_{n+1}=\left(\sum_{j=0}^{p}c_{j}^{pq}\frac{1}{j!}\Delta t^{j}D_{n,j}\right){\mathbf{w}}_{n},$ (15)
so that a single timestep consists of solving a single linear system of $N$ equations:
${\mathbf{w}}_{n+1}-L_{n+1,q}{\mathbf{w}}_{n+1}={\mathbf{w}}_{n}+R_{n,p}{\mathbf{w}}_{n},$ (16)
where
$L_{n+1,q}\coloneqq-\sum_{j=1}^{q}(-1)^{j}c_{j}^{qp}\frac{1}{j!}\Delta t^{j}D_{n+1,j},\quad R_{n,p}\coloneqq\sum_{j=1}^{p}c_{j}^{pq}\frac{1}{j!}\Delta t^{j}D_{n,j}.$ (17)
We can either form the $N\times N$ matrices $L_{n+1,q}$ and $R_{n,p}$ explicitly by explicitly forming the matrices $D_{n,j}$ and then adding them as in (17), or we compute $L_{n+1,q}{\mathbf{w}}_{n+1}$ and $R_{n,p}{\mathbf{w}}_{n}$ in a “matrix-free” fashion by applying $D_{n+1,j}$ and $D_{n,j}$ to ${\mathbf{w}}_{n+1}$ and ${\mathbf{w}}_{n}$, and adding the resulting vectors. Furthermore, by storing $D_{n+1,j}{\mathbf{w}}_{n+1}$ and $D_{n,j}{\mathbf{w}}_{n}$ after we compute them in the first $j$ terms of the summations in (15), we can compute $D_{n+1,j+1}{\mathbf{w}}_{n+1}$ and $D_{n,j+1}{\mathbf{w}}_{n}$ with only an additional $j$ matrix-vector multiplications. Therefore it costs $q(q+1)/2$ and $p(p+1)/2$ matrix-vector multiplications to compute $L_{n+1,q}{\mathbf{w}}_{n+1}$ and $R_{n+1,q}{\mathbf{w}}_{n}$, respectively, so the cost scales quadratically with the order of the method.
In this paper, we always use the matrix-free version. We store only the matrices $A_{n}^{(j)}$, which are sparse for most quantum systems, and solve (16) iteratively, using Algorithm 1 to apply $-L_{n+1,q}$ and $R_{n,p}$.
Algorithm 1 Applying $-L_{n,q}$ (or $R_{n,p})$ to ${\mathbf{w}}_{n}$.
For the system of ODEs $\frac{d{\mathbf{w}}}{d{t}}=A(t;\boldsymbol{\theta}){\mathbf{w}}$, compute $-L_{n,q}{\mathbf{w}}_{n}$ (or $R_{n,p}{\mathbf{w}}_{n}$), as required for the timestepping rule (16). When $p=q$, the timestepping method is A-stable and has order of accuracy $2p$.
$\frac{{\mathbf{w}}_{n}^{j}}{j!}\leftarrow\frac{1}{j}\sum_{i=0}^{j-1}\frac{1}{(j-1-i)!}A_{n}^{(j-1-i)}\frac{{\mathbf{w}}_{n}^{(i)}}{i!}$ for $j=1,\dots,q$.
return $\sum_{j=1}^{q}(-1)^{j}c_{j}^{qp}\Delta t^{j}\frac{{\mathbf{w}}_{n}^{(j)}}{j!}$triangleright To apply $R_{n,p}$ instead of $L_{n,q}$, swap $p$ and $q$ and remove the $(-1)^{j}$ factor.
Remark 1.
Although we do not construct the linear operators $L$ and $R$, the simple form of the timestepping (16) when written in terms of these linear operators will make deriving the discrete adjoint in Section 3.4 simple.
3.3 Control Pulse Ansatz
It should be noted for the quantum optimal control problem that in addition to the matrix $A$ being the sum of a drift and control matrix, as in (3), the time-dependence takes the specific form
$A(t)=A_{d}+A_{{\mathpzc c}}(t),\ \ \text{where}\quad A_{\mathpzc c}(t)=\sum_{j=1}^{N_{C}}{\mathpzc c}_{j}(t)A_{{\mathpzc c},j},\ \quad{\mathpzc c}_{j}(t)=\sum_{k=1}^{n_{j}}\theta_{j,k}{\mathpzc c}_{j,k}(t),\quad\sum_{j=1}^{{N_{\rm C}}}n_{j}={N_{\rm P}}.$ (18)
Here the multi-index $\theta_{j,k}$ simply indicates the $k$-th control parameter for the $j$-th control pulse (it is also assumed here that each control parameter affects only one control pulse, though in principle this need not be the case). Each ${\mathpzc c}_{j}(t)$ is a scalar function expressed as a linear combination of pre-defined control functions which we call the control pulse ansatz. A nonlinear dependence could also be considered, but constructing the control pulses this way makes it easy to compute their time derivatives and their gradients with respect to the control parameters. It can be done in the following way:
$\frac{d^{i}{\mathpzc c}_{j}(t)}{d{t}^{i}}=\sum_{k=1}^{n_{j}}\theta_{j,k}\frac{d^{i}{\mathpzc c}_{j,k}(t)}{d{t}^{i}},\quad\frac{\partial}{\partial\theta_{j,k}}\frac{d^{i}{\mathpzc c}_{j}(t)}{d{t}^{i}}=\frac{d^{i}{\mathpzc c}_{j,k}(t)}{d{t}^{i}},$
and we can choose basis functions ${\mathpzc c}_{j,k}$ whose time derivatives can be easily computed.
The only requirements with respect to the Hermite timestepping method are that we may compute the value of the control pulse and its time derivatives, as well as their gradients with respect to the control parameters, and that the control pulses are sufficiently smooth to match the order of accuracy of the method being used. In principle, any control pulse ansatz may be used.
As our particular choice of control pulse ansatz, we follow the approach of [2] and use B-splines (basis splines) [8] multiplied by carrier waves, noting the importance of good frequency selection in achieving low infidelities with a small number of control parameters [6]. B-splines have several properties that make them suitable for use as a control pulse ansatz. 1) The partition of unity property ensures that the maximum value of a B-spline curve does not exceed the maximum control coefficient of the basis splines, therefore a maximum amplitude of the control pulses can be limited by placing a maximum absolute value on the control coefficients as a constraint in the optimization problem. 2) The local support property makes the pulse parametrization intuitive, since each control coefficient only affects the B-spline curve in a finite, compact region. This also makes it easy to design control pulses which begin and end with amplitude zero, which can be experimentally desirable. 3) The number of continuous derivatives of the B-spline curve increases with the order of basis splines used. Consequently, we can build arbitrarily smooth control pulses by using higher-order basis splines. 4) Finally, there are established, robust, and efficient software implementations of the evaluation of basis splines and B-spline curves [8].
3.4 Discrete Adjoint
A discrete adjoint method for evaluating derivatives of functions of discrete sequences implicitly defined through forward difference equations has been given by Betancourt et al. [3]. We extend that treatment here to be compatible with our method. Let the objective function be defined by the discrete functional
$\mathcal{J}(\boldsymbol{\theta})=\sum_{n=0}^{{N_{\rm T}}}{\mathfrak{j}}_{n}({\mathbf{w}}_{n},\boldsymbol{\theta}).$
Here we describe the case where the state is a vector (e.g. a state preparation problem), but the result can easily be extended to the matrix (gate design) case.
The derivatives of $\mathcal{J}$, which we want to compute in order to perform a gradient-based search, are given by
$\frac{d\mathcal{J}}{d{\theta_{k}}}=\sum_{n=0}^{N}\frac{\partial{\mathfrak{j}}_{n}}{\partial\theta_{k}}+\left(\frac{\partial{\mathfrak{j}}_{n}}{\partial{\mathbf{w}}_{n}}\right)^{T}\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}}.$ (19)
The ${\partial{\mathfrak{j}}_{n}}/{\partial\theta_{k}}$ terms are zero unless the objective function depends explicitly on $\boldsymbol{\theta}$, e.g. if it includes a regularization term $\|\boldsymbol{\theta}\|_{2}$. For simplicity, we will assume that each ${\mathfrak{j}}_{n}$ only depends $\boldsymbol{\theta}$ implicitly through the dependence of ${\mathbf{w}}_{n}$ on $\boldsymbol{\theta}$. The trajectory of ${\mathbf{w}}$ is given by our timestepping rule:
${\mathbf{w}}_{n+1}-L_{n+1,q}{\mathbf{w}}_{n+1}={\mathbf{w}}_{n}+R_{n,p}{\mathbf{w}}_{n},\quad n=0,1,\dots,{N_{\rm T}}-1.$ (20)
Computing $d\mathcal{J}/d\boldsymbol{\theta}$ directly using (19) requires computing the sensitivities $d{\mathbf{w}}_{n}/d\theta_{k}$ for each control parameter, which requires us to evolve a Schrödinger equation for each control parameter. Instead, we treat the method as a constraint in the method of Lagrange multipliers, giving us the Lagrangian
$\mathcal{L}(\boldsymbol{\theta})$ $=\mathcal{J}(\boldsymbol{\theta})+\sum_{n=0}^{{N_{\rm T}}-1}\left\langle{\mathbf{w}}_{n+1}-L_{n+1,q}{\mathbf{w}}_{n+1}-{\mathbf{w}}_{n}-R_{n,p}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$ (21)
$={\mathfrak{j}}_{N_{\rm T}}+\sum_{n=0}^{{N_{\rm T}}-1}{\mathfrak{j}}_{n}+\left\langle{\mathbf{w}}_{n+1}-L_{n+1,q}{\mathbf{w}}_{n+1}-{\mathbf{w}}_{n}-R_{n,p}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle.$ (22)
If the constraints (20) are satisfied, then $\mathcal{L}(\boldsymbol{\theta})=\mathcal{J}(\boldsymbol{\theta})$ for any value of the Lagrange multipliers $\boldsymbol{\lambda}_{n}$. Because the constraints are defined using the exact timestepping scheme we use to compute each ${\mathbf{w}}_{n}$, we may assume the constraints are always satisfied and therefore $\mathcal{L}(\boldsymbol{\theta})\equiv\mathcal{J}(\boldsymbol{\theta})$. Taking $\frac{d}{d{\theta_{k}}}$ and making use of the fact that each $\boldsymbol{\lambda}_{n}$ has no explicit dependence on $\boldsymbol{\theta}$ (we are free to choose $\boldsymbol{\lambda}_{n}$ as we like), we get
$\frac{d\mathcal{L}}{d{\theta_{k}}}=\frac{d\mathcal{J}}{d{\theta_{k}}}=\frac{d{\mathfrak{j}}_{N_{\rm T}}}{d{\theta_{k}}}+\sum_{n=0}^{{N_{\rm T}}-1}\frac{d{\mathfrak{j}}_{n}}{d{\theta_{k}}}+\left\langle\frac{d{\mathbf{w}}_{n+1}}{d{\theta_{k}}}-\frac{d}{d{\theta_{k}}}\left(L_{n+1,q}{\mathbf{w}}_{n+1}\right)-\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}}-\frac{d}{d{\theta_{k}}}\left(R_{n,p}{\mathbf{w}}_{n}\right),\boldsymbol{\lambda}_{n+1}\right\rangle.$ (23)
Our goal is to rewrite (23) and choose $\lambda_{n}$ such that they cancel out the sensitivities $d{\mathbf{w}}_{n}/d\theta_{k}$, leaving us with a gradient which is much easier to compute.
Theorem 1.
Let the numerical approximation to Schrödinger’s equation (1) satisfy the timestepping rule (20). Then the exact gradient of the discrete cost function (21) with respect to the numerical method (12) can be computed by
$\frac{d\mathcal{J}}{d{\theta_{k}}}=-\sum_{n=0}^{{N_{\rm T}}-1}\left\langle\frac{\partial L_{n+1,q}}{\partial\theta_{k}}{\mathbf{w}}_{n+1}+\frac{\partial R_{n,p}}{\partial\theta_{k}}{\mathbf{w}}_{n},\ \boldsymbol{\lambda}_{n+1}\right\rangle,$ (24)
where $\boldsymbol{\lambda}_{n}$ is chosen to satisfy the terminal condition and adjoint equations [^7] We follow the convention that the partial derivative of a scalar with respect to a column vector is a row vector.
$\boldsymbol{\lambda}_{N_{\rm T}}-(L_{{N_{\rm T}},q})^{T}\boldsymbol{\lambda}_{N_{\rm T}}$ $=-\left(\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\right)^{T},$ (25)
$\boldsymbol{\lambda}_{n}-(L_{n,q})^{T}\boldsymbol{\lambda}_{n}$ $=\boldsymbol{\lambda}_{n+1}+(R_{n,p})^{T}\boldsymbol{\lambda}_{n+1}-\left(\frac{\partial{\mathfrak{j}}_{n}}{\partial{\mathbf{w}}_{n}}\right)^{T},\quad(n={N_{\rm T}}-1,\dots,1).$ (26)
To this end we refer to the solving of the linear systems in (25) and (26) as the adjoint evolution and the explicit computation of the gradient in equation (24) as the gradient accumulation. The entire procedure of performing the forward evolution, adjoint evolution, and gradient accumulation is summarized in Algorithm 2.
Proof.
Applying the chain rule to (23) yields
$\frac{d\mathcal{J}}{d{\theta_{k}}}=\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\frac{d{\mathbf{w}}_{N_{\rm T}}}{d{\theta_{k}}}+\sum_{n=0}^{{N_{\rm T}}-1}\frac{\partial{\mathfrak{j}}_{n}}{\partial{\mathbf{w}}_{n}}\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}}\\ +\sum_{n=0}^{{N_{\rm T}}-1}\left\langle\frac{d{\mathbf{w}}_{n+1}}{d{\theta_{k}}}-\frac{\partial\left(L_{n+1,q}{\mathbf{w}}_{n+1}\right)}{\partial\theta_{k}}-\frac{\partial\left(L_{n+1,q}{\mathbf{w}}_{n+1}\right)}{\partial{\mathbf{w}}_{n+1}}\frac{d{\mathbf{w}}_{n+1}}{d{\theta_{k}}}-\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}}-\frac{\partial\left(R_{n,p}{\mathbf{w}}_{n}\right)}{\partial\theta_{k}}-\frac{\partial\left(R_{n,p}{\mathbf{w}}_{n}\right)}{\partial{\mathbf{w}}_{n}}\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}},\boldsymbol{\lambda}_{n+1}\right\rangle.$
For a matrix vector product $M\mathbf{x}$, we have $\partial(M\mathbf{x})/\partial\mathbf{x}=M$. Also, the state vectors have no explicit dependence on $\boldsymbol{\theta}$ (the dependence is only implicit), so the above expression for ${d\mathcal{J}}/{d{\theta_{k}}}$ simplifies to
$\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\frac{d{\mathbf{w}}_{N_{\rm T}}}{d{\theta_{k}}}\\ +\sum_{n=0}^{{N_{\rm T}}-1}\frac{\partial{\mathfrak{j}}_{n}}{\partial{\mathbf{w}}_{n}}\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}}+\left\langle\frac{d{\mathbf{w}}_{n+1}}{d{\theta_{k}}}-\frac{\partial L_{n+1,q}}{\partial\theta_{k}}{\mathbf{w}}_{n+1}-L_{n+1,q}\frac{d{\mathbf{w}}_{n+1}}{d{\theta_{k}}}-\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}}-\frac{\partial R_{n,p}}{\partial\theta_{k}}{\mathbf{w}}_{n}-R_{n,p}\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}},\boldsymbol{\lambda}_{n+1}\right\rangle.$
We rearrange the terms to make it clear how to choose $\boldsymbol{\lambda}$ to cancel out the sensitivities:
$\frac{d\mathcal{J}}{d{\theta_{k}}}=\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\frac{d{\mathbf{w}}_{N_{\rm T}}}{d{\theta_{k}}}+\left\langle\frac{d{\mathbf{w}}_{N_{\rm T}}}{d{\theta_{k}}}-L_{{N_{\rm T}},q}\frac{d{\mathbf{w}}_{{N_{\rm T}}}}{d{\theta_{k}}},\boldsymbol{\lambda}_{N_{\rm T}}\right\rangle\\ +\sum_{n=0}^{{N_{\rm T}}-1}\frac{\partial{\mathfrak{j}}_{n}}{\partial{\mathbf{w}}_{n}}\frac{d{\mathbf{w}}_{k}}{d{\theta_{k}}}+\left\langle\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}}-L_{n,q}\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}},\boldsymbol{\lambda}_{n}\right\rangle-\left\langle\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}}+R_{n,p}\frac{d{\mathbf{w}}_{n}}{d{\theta_{k}}},\boldsymbol{\lambda}_{n+1}\right\rangle-\left\langle\frac{\partial L_{n+1,q}}{\partial\theta_{k}}{\mathbf{w}}_{n+1}+\frac{\partial R_{n,p}}{\partial\theta_{k}}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle.$
Now, making use of the identities $\left\langle\boldsymbol{a}\boldsymbol{b},\boldsymbol{c}\right\rangle=\left\langle\boldsymbol{b},\boldsymbol{a}^{T}\boldsymbol{c}\right\rangle$ and $\boldsymbol{a}^{T}\boldsymbol{b}=\boldsymbol{b}^{T}\boldsymbol{a}$, we get
$\frac{d\mathcal{J}}{d{\theta_{k}}}=\left\langle\frac{d{\mathbf{w}}_{N_{\rm T}}}{d{\theta_{k}}},\left(\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\right)^{T}+\boldsymbol{\lambda}_{N_{\rm T}}-\left(L_{{N_{\rm T}},q}\right)^{T}\boldsymbol{\lambda}_{N_{\rm T}}\right\rangle\\ +\sum_{n=0}^{{N_{\rm T}}-1}\left\langle\frac{d{\mathbf{w}}_{k}}{d{\theta_{k}}},\left(\frac{\partial{\mathfrak{j}}_{n}}{\partial{\mathbf{w}}_{n}}\right)^{T}+\boldsymbol{\lambda}_{n}-\left(L_{n,q}\right)^{T}\boldsymbol{\lambda}_{n}-\boldsymbol{\lambda}_{n+1}-\left(R_{n,p}\right)^{T}\boldsymbol{\lambda}_{n+1}\right\rangle-\left\langle\frac{\partial L_{n+1,q}}{\partial\theta_{k}}{\mathbf{w}}_{n+1}+\frac{\partial R_{n,p}}{\partial\theta_{k}}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle.$ (27)
If we choose $\boldsymbol{\lambda}$ to satisfy the terminal condition and adjoint equations (25) and (26), then we will have
$\left(\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\right)^{T}+\boldsymbol{\lambda}_{N_{\rm T}}-(L_{{N_{\rm T}},q})^{T}\boldsymbol{\lambda}_{N_{\rm T}}$ $=\boldsymbol{0}$
$\left(\frac{\partial{\mathfrak{j}}_{n}}{\partial{\mathbf{w}}_{n}}\right)^{T}+\boldsymbol{\lambda}_{n}-(L_{n,q})^{T}\boldsymbol{\lambda}_{n}-\boldsymbol{\lambda}_{n+1}-(R_{n,p})^{T}\boldsymbol{\lambda}_{n+1}$ $=\boldsymbol{0},\quad(n=0,\dots,{N_{\rm T}}-1).$
Consequently all occurrences of $d{\mathbf{w}}_{n}/d\theta_{k}$ in $d\mathcal{J}/d\theta_{k}$ as written in (27) will have their inner product taken with a quantity equal to $\boldsymbol{0}$, so that the only remaining nonzero terms are exactly those given in (24). ∎
Algorithm 2 The HOHO Method / Discrete Adjoint
Given the initial value problem $\frac{d{\mathbf{w}}}{d{t}}=A(t;\boldsymbol{\theta}){\mathbf{w}},\quad 0\leq t\leq T,\quad{\mathbf{w}}(0)={\mathbf{w}}_{0},$ and the discrete objective functional $\mathcal{J}=\sum_{n=1}^{{N_{\rm T}}}{\mathfrak{j}}({\mathbf{w}}_{n})$, efficiently compute the gradient $\nabla\mathcal{J}=[{d\mathcal{J}}/{d{\theta_{1}}},\dots,{d\mathcal{J}}/{d{\theta_{N_{\rm C}}}}]^{T}$ exactly with respect to the Hermite Runge-Kutta method using $q$ and $p$ derivatives for the implicit/explicit parts.
Forward Evolution (Solve Forward Equations)$\triangleright$ See Algorithm 1.
for $n=1:{N_{\rm T}}$ do
     Compute ${\mathbf{w}}_{n}+R_{n,p}{\mathbf{w}}_{n}$ explicitly.
     Solve ${\mathbf{w}}_{n+1}-L_{n+1,q}{\mathbf{w}}_{n+1}={\mathbf{w}}_{n}+R_{n,p}{\mathbf{w}}_{n}$ for ${\mathbf{w}}_{n+1}$ using the most efficient method available.
end for
Adjoint Evolution (Solve Terminal Condition and Adjoint Equations)$\triangleright$ See Algorithm 4.
Compute $-\left(\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\right)$ explicitly.
Solve $\boldsymbol{\lambda}_{N_{\rm T}}-\left(L_{{N_{\rm T}},q}\right)^{T}\boldsymbol{\lambda}_{N_{\rm T}}=-\left(\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\right)$ for $\boldsymbol{\lambda}_{N_{\rm T}}$ using the most efficient method available.
for $n={N_{\rm T}}-1:-1:1$ do
     Compute $\boldsymbol{\lambda}_{n+1}+\left(R_{n,p}\right)^{T}\boldsymbol{\lambda}_{n+1}-\left(\frac{\partial{\mathfrak{j}}_{n}}{\partial{\mathbf{w}}_{n}}\right)$ explicitly.
     Solve $\boldsymbol{\lambda}_{n}-\left(L_{n,q}\right)^{T}\boldsymbol{\lambda}_{n}=\boldsymbol{\lambda}_{n+1}+\left(R_{n,p}\right)^{T}\boldsymbol{\lambda}_{n+1}-\left(\frac{\partial{\mathfrak{j}}_{N_{\rm T}}}{\partial{\mathbf{w}}_{N_{\rm T}}}\right)$ for $\boldsymbol{\lambda}_{n}$ using the most efficient method available.
end for
[…]
$\nabla\mathcal{J}\leftarrow\boldsymbol{0}$
$(\nabla\mathcal{J})_{k}\leftarrow(\nabla\mathcal{J})_{k}-\left\langle\frac{\partial L_{n+1,q}}{\partial\theta_{k}}{\mathbf{w}}_{n+1}+\frac{\partial R_{n,p}}{\partial\theta_{k}}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$ for $n=1,\dots,{N_{\rm T}},\quad k=1,\dots,{N_{\rm P}}$
return $\nabla\mathcal{J}$
Remark 2.
When the state is a matrix (e.g. a gate design problem), the gradient may be computed by performing the forward evolution, adjoint evolution, and gradient accumulation for each initial condition and summing the results. The computation is almost completely independent for each initial condition, except that computing ${\partial{\mathfrak{j}}_{n}}/{\partial{\mathbf{w}}_{n}}$, where ${\mathbf{w}}_{n}$ is the state vector after $n$ timesteps of a single intial condition may require the state vector after the $n$h timestep for every initial condition. Typically, this is only the case for the final time objective ${\mathfrak{j}}_{N_{\rm T}}$ (e.g. the infidelity). Therefore, we can perfectly parallelize the forward evolution across the intial conditions, then solve the terminal condition equation (25), then perform the rest of the adjoint evolution (26) and gradient accumulation (24) in parallel.
Unlike the forward and adjoint evolutions, performing the gradient accumulation (24) does not require solving any linear systems; it only requires explicit matrix-vector multiplications. But when the forward and adjoint evolutions are performed using a small timestep size $\Delta t$, a good initial guess for each timestep, and/or an effective preconditioner, the linear systems may be solved in only a few iterations, requiring only a few matrix-vector multiplications. In that case, because the number of matrix-vector multiplications required by the gradient accumulation as written in (24) scales with the number of control parameters and the number of timesteps, the gradient accumulation can be much more expensive than the forward and adjoint evolutions if handled naively. We now show how to perform the gradient accumulation (24) using a number of matrix-vector multiplications that is independent of the number of control parameters.
Theorem 2.
When the time-dependence of $A(t)$ has the structure (18), the gradient accumulation (24) can be rewritten so that the number of matrix-vector multiplications required is independent of the number of control parameters ${N_{\rm C}}$.
Proof.
Expanding $L_{n+1,q}$ and $R_{n,p}$ in (24) (and using the fact that $D_{n,0}=I$ does not depend on $\boldsymbol{\theta}$), we get
$\frac{d\mathcal{J}}{d{\theta_{k}}}$ $=-\sum_{n=0}^{{N_{\rm T}}-1}\left(\left\langle-\sum_{j=1}^{q}(-1)^{j}c_{j}^{qp}\frac{1}{j!}\frac{\partial D_{n+1,j}}{\partial\theta_{k}}{\mathbf{w}}_{n+1}+\sum_{j=1}^{p}c_{i}^{pq}\frac{1}{j!}\frac{\partial D_{n,j}}{\partial\theta_{k}}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle\right)$ (28)
$=\sum_{n=0}^{{N_{\rm T}}-1}\left(\sum_{j=1}^{q}(-1)^{j}c_{j}^{qp}\frac{1}{j!}\left\langle\frac{\partial D_{n+1,j}}{\partial\theta_{k}}{\mathbf{w}}_{n+1},\boldsymbol{\lambda}_{n+1}\right\rangle-\sum_{j=1}^{p}c_{j}^{pq}\frac{1}{j!}\left\langle\frac{\partial D_{n,j}}{\partial\theta_{k}}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle\right).$ (29)
To compute the inner products efficiently, we take $\partial/\partial\theta_{k}$ of (14):
$\frac{\partial D_{n,0}}{\partial\theta_{k}}=0,\quad\frac{\partial D_{n,j+1}}{\partial\theta_{k}}=\sum_{i=0}^{j}\binom{j}{i}\frac{\partial A_{n}^{(j-i)}}{\partial\theta_{k}}D_{n,i}+A_{n}^{(j-i)}\frac{\partial D_{n,i}}{\partial\theta_{k}},\quad j=0,\dots,q-1,$
and use the identity $\left\langle M{\mathbf{a}},{\mathbf{b}}\right\rangle=\left\langle{\mathbf{a}},M^{T}{\mathbf{b}}\right\rangle$ to rewrite the inner product as
$\left\langle\frac{\partial D_{n,j+1}}{\partial\theta_{k}}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$ $=\sum_{i=0}^{j}\binom{j}{i}\left\langle\frac{\partial A_{n}^{(j-i)}}{\partial\theta_{k}}D_{n,i}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle+\binom{j}{i}\left\langle A_{n}^{(j-i)}\frac{\partial D_{n,i}}{\partial\theta_{k}}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$ (30)
$=\sum_{i=0}^{j}\binom{j}{i}\left\langle\frac{\partial A_{n}^{(j-i)}}{\partial\theta_{k}}D_{n,i}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle+\binom{j}{i}\left\langle\frac{\partial D_{n,i}}{\partial\theta_{k}}{\mathbf{w}}_{n},\left(A_{n}^{(j-i)}\right)^{T}\boldsymbol{\lambda}_{n+1}\right\rangle.$ (31)
The inner products contatining $\partial A_{n}^{(j-i)}/\partial\theta_{k}$ in (31) can be handled explicitly (and more efficiently by also storing each $D_{n,i}{\mathbf{w}}_{n}$ to use when computing $D_{n,i+1}{\mathbf{w}}_{n}$, as described in Section 3.2), while the inner products containing $\partial D_{n,i}/\partial\theta_{k}$ can be handled recursively by expanding the inner product using (30). Ultimately, the only matrix-vector multiplications performed will be in the computation of the first argument of inner products of the form $\left\langle\frac{\partial A_{n}^{(i)}}{\partial\theta_{k}}\mathbf{x},{\mathbf{y}}\right\rangle$.
The number of matrix-vector multiplications scales with the number of control parameters because the matrix ${\partial A_{n}^{(i)}}/{\partial\theta_{k}}$ may be different for each $\theta_{k}$. However, when $A(t)$ has the form (18) the inner product may be rewritten as
$\left\langle\frac{\partial A_{n}^{(i)}}{\partial\theta_{k}}\mathbf{x},{\mathbf{y}}\right\rangle=\sum_{j=1}^{N_{\rm C}}\frac{\partial c_{j}^{(i)}(t_{n})}{\partial\theta_{k}}\left\langle A_{{\mathpzc c},j}\mathbf{x},{\mathbf{y}}\right\rangle=\sum_{j=1}^{N_{\rm C}}\frac{\partial c_{j}^{(i)}(t_{n})}{\partial\theta_{k}}\left\langle\mathbf{x},A^{T}_{{\mathpzc c},j}{\mathbf{y}}\right\rangle,$ (32)
which allows us to compute the inner product by performing only ${N_{\rm C}}$ (the number of controls) matrix-vector multiplications, independent of the number of control parameters, completing our proof. ∎
Remark 3.
Moving $A_{{\mathpzc c},j}$ to the second argument of the inner product in (32) allows us to compute the two inner products in (29) using the same matrix-vector multiplications, since the inner products have the same second argument. I.e., the gradient accumulation requires us to compute inner products
$\left\langle\frac{\partial A_{n}^{(i)}}{\partial\theta_{k}}\mathbf{x}_{1},{\mathbf{y}}\right\rangle=\sum_{j=1}^{N_{\rm C}}\frac{\partial c_{j}^{(i)}(t_{n})}{\partial\theta_{k}}\left\langle\mathbf{x}_{1},A^{T}_{{\mathpzc c},j}{\mathbf{y}}\right\rangle,\quad\left\langle\frac{\partial A_{n+1}^{(i)}}{\partial\theta_{k}}\mathbf{x}_{2},{\mathbf{y}}\right\rangle=\sum_{j=1}^{N_{\rm C}}\frac{\partial c_{j}^{(i)}(t_{n+1})}{\partial\theta_{k}}\left\langle\mathbf{x}_{2},A^{T}_{{\mathpzc c},j}{\mathbf{y}}\right\rangle,$
and by reusing $A_{c,j}^{T}{\mathbf{y}}$ we can compute both inner products with only ${N_{\rm C}}$ matrix-vector multiplications. This halves the number of matrix-vector multiplications required for the gradient accumulation.
The efficient gradient accumulation techniques detailed in Theorem 2 and Remark 3 can by applied by using Algorithm 3.
Algorithm 3 Efficient Gradient Accumulation
Compute $\nabla_{\boldsymbol{\theta}}\left\langle L_{n+1,q}{\mathbf{w}}_{n+1}+R_{n,p}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$, in order to perform the gradient accumulation (24) efficiently when $A$ takes the usual form: $A(t;\boldsymbol{\theta})=A_{0}+c_{1}(t;\boldsymbol{\theta})A_{1}+\dots+c_{N_{\rm C}}(t;\boldsymbol{\theta})A_{N_{\rm C}}$. Assumes $p=q$, so that the recursive procedures for computing $\nabla_{\boldsymbol{\theta}}\left\langle L_{n+1,q}{\mathbf{w}}_{n+1},\boldsymbol{\lambda}_{n+1}\right\rangle$ and $\nabla_{\boldsymbol{\theta}}\left\langle R_{n,q}{\mathbf{w}}_{n},\boldsymbol{\lambda}_{n+1}\right\rangle$ have the same depth.
Precompute $A_{k}\frac{{\mathbf{w}}_{n+1}^{(i)}}{i!}$ and $A_{k}\frac{{\mathbf{w}}_{n}^{(i)}}{i!}$ for $i=0,\dots,q$, $k=1,\dots,{N_{\rm C}}$.
return $\sum_{j=1}^{q}\textsc{$\nabla$D}(\boldsymbol{\lambda}_{n+1},\ j,\ (-1)^{j+1}c_{j}^{qp}\Delta t^{j},\ c_{j}^{pq}\Delta t^{j})$
[…]
     return $\frac{1}{j}\sum_{i=0}^{j-1}\left(\nabla\textrm{D}(\frac{1}{(j-1-i)!}A^{(j-1-i)}\boldsymbol{\lambda},\ i,\ c_{L},\ c_{R})+\sum_{k=1}^{N_{\rm C}}\frac{\nabla{\mathpzc c}_{n+1,k}^{(j-1-i)}}{(j-1-i)!}\left\langle A_{k}\frac{{\mathbf{w}}_{n+1}^{(i)}}{i!},\boldsymbol{\lambda}\right\rangle+\frac{\nabla{\mathpzc c}_{n,k}^{(j-1-i)}}{(j-1-i)!}\left\langle A_{k}\frac{{\mathbf{w}}_{n}^{(i)}}{i!},\boldsymbol{\lambda}\right\rangle\right)$
end procedure
Remark 4.
Applying $(L_{n,q})^{T}$ and $(R_{n,p})^{T}$ when solving the adjoint equations requires more matrix-vector multiplications than applying $L_{n,q}$ and $R_{n,p}$ in the forward evolution.
Taking the transpose of (14), we get
$(D_{n,0})^{T}=I,\quad(D_{n,j+1})^{T}=\sum_{i=0}^{j}\binom{j}{i}(D_{n,i})^{T}\left(A_{n}^{(j-i)}\right)^{T},\quad j=0,\dots,q-1.$ (33)
Whereas the computation of $D_{n,j+1}\mathbf{x}$ can be done efficiently by storing and reusing $D_{n,i}\mathbf{x}$ for $i=1,\dots,j$, in the computation of $(D_{n,j+1})^{T}\mathbf{x}$ the vector being multiplied by each $(D_{n,i})^{T}$ changes with the values of $i$ and $j$. The vectors $A_{n}^{(i)}\mathbf{x}$, $i=0,\dots,j$ may be computed once and and reused. When this is done, the number of matrix-vector mutiplications needed to multiply a vector by $(D_{n,j+1})^{T}$ is
${\rm cost}[(D_{n,j+1})^{T}]=1+\sum_{i=0}^{j}{\rm cost}[(D_{n,j})^{T}],\quad j=1,\dots,q,$
and the number of matrix-vector multiplications needed to multiply a vector by $L_{n,q}$ or $R_{n,p}$ is
${\rm cost}[(L_{n,q})^{T}]=\sum_{i=0}^{q-1}{\rm cost}[(D_{n,j})^{T}],\quad{\rm cost}[(R_{n,p})^{T}]=\sum_{i=0}^{p-1}{\rm cost}[(D_{n,j})^{T}].$
Algorithm 4 gives a recursive method for applying $(L_{n,q})^{T}$ and $(R_{n,p})^{T}$ when solving the adjoint equations.
A comparison of the number of matrix-vector multiplications required in the application of $L_{n,q}$ or $R_{n,p}$ (as in the forward evolution) and in the application of $(L_{n,q})^{T}$ or $(R_{n,p})^{T}$ (as in the adjoint evolution) when $q=p$ is given in Table 1. The number of matrix-vector multiplications required in the adjoint evolution scales exponentially with the order of the method, limiting the efficiency of the extremely high order methods. In the rest of this work, we consider method orders twelve and under, for which the number of matrix-vector multiplications required in the adjoint evolution does not exceed three times the number required in the forward evolution.
Order of Method
2 4 6 8 10 12 14 16
# Matvec Mults: Forward 1 3 6 10 15 21 28 36
# Matvec Mults: Adjoint 1 3 7 15 31 63 127 255
Table 1: Number of matrix-vector multiplications required in the application of $L_{n,q}$ and $R_{n,p}$ (as in the forward evolution) and in the application of $(L_{n,q})^{T}$ and $(R_{n,p})^{T}$ (as in the adjoint evolution) when $q=p$.
Algorithm 4 Applying $-(L_{n,q})^{T}$ (or $(R_{n,p})^{T}$) to $\boldsymbol{\lambda}_{n}$.
For the system of ODEs $\frac{d{\mathbf{w}}}{d{t}}=A(t;\boldsymbol{\theta}){\mathbf{w}}$, given $\boldsymbol{\lambda}_{n}$, the adjoint state corresponding to the $n$-th timestep, compute $-(L_{n,q})^{T}\boldsymbol{\lambda}_{n}$ (or $(R_{n,p})^{T}\boldsymbol{\lambda}_{n}$), as required when solving the teriminal condition (25) or adjoint equations (26).
Precompute $\ \frac{1}{(j-1)!}(A_{n}^{(j-1)})^{T}\boldsymbol{\lambda}_{n}$ for $j=1,\dots,q$.
return $\sum_{j=1}^{q}(-1)^{j}c_{j}^{qp}\Delta t^{j}\textsc{DT}(\boldsymbol{\lambda}_{n},j)$triangleright To apply $R_{n,p}$ instead of $L_{n,q}$, swap $p$ and $q$ and remove the $(-1)^{j}$ factor.
procedure DT($\mathbf{x}$, $j$) $\longrightarrow\frac{1}{j!}(D_{n,j})^{T}\mathbf{x}$
     if $j=0$ then
         return $\mathbf{x}$
     $\triangleright$ (When DT is not called recursively, then $\mathbf{x}=\boldsymbol{\lambda}_{n}$ and $\frac{1}{(j-i)!}(A_{n}^{(j-i)})^{T}\mathbf{x}$ is precomputed.)
         return $\frac{1}{j}\sum_{i=0}^{j-1}\textsc{DT}(\frac{1}{(j-i)!}(A_{n}^{(j-i)})^{T}\mathbf{x},i)$
     end if
end procedure
Remark 5.
Because the terminal condition (25) and adjoint equations (26) require computing $({\partial{\mathfrak{j}}_{n}}/{\partial{\mathbf{w}}_{n}})^{T}$ for $n={N_{\rm T}},\dots,1$, it is necessary to either store the entire state history ${\mathbf{w}}_{0},\dots,{\mathbf{w}}_{N_{\rm T}}$ when solving the forward equations, or to solve Schrödinger’s equation backward in time while solving the adjoint equations, so that we only every store a single state vector or adjoint state vector at each timestep. We refer to this as the “memory-lean” version. In this case, the gradient accumulation (24) is also performed alongside the adjoint evolution.
There are two main drawbacks to the memory-lean version. 1) We must solve an additional system of ODEs, increasing the cost of our method by roughly 50%. 2) Because the Hermite Runge-Kutta methods we use are non-reversible, the memory-lean version computed gradient will no longer be exact with respect to the discretized forward equations. The latter drawback can be especially detrimental to the performance of quasi-Newton optimizers. As a compromise, a checkpointing approach [15] could be used to store the state vector at only a subset of of all timesteps and still computing an accurate gradient with respect to the discretized forward equations.
3.5 Preconditioning
We find that an efficient left preconditioner for the solution of the linear system (16) is to apply $(I-\tilde{L}_{n+1,q})^{-1}$, where $\tilde{L}_{n+1,q}$ is the operator $L_{n+1,q}$ in the case where the time-dependent Hamiltonian $A(t)$ is replaced with the time-independent drift Hamiltonian $A_{d}$ (i.e. $A(t)$ when the control pulse amplitudes are all zero). Typically, the control pulse amplitudes are experimentally required to remain below a value which is significantly smaller than the eigenvalues of the drift Hamiltonian. Combined with the fact that $L_{n+1,q}$ scales with $\Delta t$, this suggests $L_{n+1,q}$ is only a small perturbation of $\tilde{L}_{n+1,q}$ for reasonably small $\Delta t$. When the drift Hamiltonian is diagonal, $(I-\tilde{L}_{n+1,q})^{-1}$ can be applied in $\mathcal{O}(N)$ time using Gaussian elimination, without ever forming $(I-\tilde{L}_{n+1,q})^{-1}$ explicitly. In this work, we consider only problems with diagonal drift Hamiltonians, and use the Gaussian elimination approach to preconditioning.
4 Numerical Experiments
4.1 Correctness Experiments on a Rabi Oscillator Problem
To verify the correctness of our numerical method, we simulate a Rabi oscillator problem, which has an analytic solution. A Rabi oscillator problem consists of a two-level system in a rotating frame of reference. For this problem, Schrödinger’s equation is
$\frac{d\boldsymbol{\psi}}{d{t}}=-iH_{c}\boldsymbol{\psi},\quad\text{where }H_{c}=\Omega a+\overline{\Omega}a^{\dagger}={\rm Re}\,\Omega(a+a^{\dagger})+{\rm Im}\,\Omega(a-a^{\dagger}),\quad\Omega\in\mathbb{C}.$ (34)
The analytic solution to (34) is $\boldsymbol{\psi}(t)=U(t)\boldsymbol{\psi}_{0}$, where the time-evolution unitary is given by
$U(t)=\begin{bmatrix}\cos(|\Omega|t)&(\sin(\theta)-i\cos(\theta))\sin(|\Omega|t)\\ -(\sin(\theta)+i\cos(\theta))\sin(|\Omega|t)&\cos(|\Omega|t)\end{bmatrix},\quad\text{where }\Omega=|\Omega|(\cos(\theta)+i\sin(\theta)).$ (35)
This solution can be derived using an eigenvector decomposition of the Hamiltonian [31, Sec. 7.23]. The populations of the $|0\rangle$ and $|1\rangle$ states oscillate between 0 and 1 with period $\tau=\pi/|\Omega|$, and hence an application of control pulses $p_{0}={\rm Re}\,(\Omega)$ and $q_{0}={\rm Im}\,(\Omega)$ for duration $\tau$ is called a Rabi oscillation. A control pulse with $p_{0}=|\Omega|$, $q_{0}=0$, and duration $\tau/2$ implements a Pauli-$X$ gate.
4.1.1 Verification of Order of Accuracy
Using the analytic solution (35), we may evaluate the error in a numerical solution of the Rabi oscillator problem (34) at any point in time for any initial condition. In Table 2, we plot the Frobenius norm of the relative error in the final-time numerical simulation of the Rabi oscillator problem for a basis of initial states $U_{0}=\operatorname{diag}(1,1)$, with $\Omega=0.05(\cos(\pi/4)+i\sin(\pi/4))$ (constant control pulses with amplitude $25/\sqrt{2}$ MHz), and $T=9.5\pi/|\Omega|$ (nine and a half Rabi oscillations), starting with sixteen timesteps and then doubling the number of timesteps until we reach 256 timesteps. We observe that the method converges to the analytic solution with the expected orders of accuracy. Furthermore, we can simulate the problem with high accuracy using far fewer timesteps with the high-order methods than with the second-order method.
Order 2 Order 4 Order 6 Order 8 Order 10 Order 12
Steps Err Cvg Err Cvg Err Cvg Err Cvg Err Cvg Err Cvg
16 4.5(-1) - 4.0(-1) - 1.1(-2) - 1.6(-4) - 1.4(-6) - 8.6(-9) -
32 1.6 1.9 3.0(-2) 3.7 1.9(-4) 5.8 6.6(-7) 7.9 1.4(-9) 9.9 2.2(-12) 11.9
64 5.2(-1) 1.7 1.9(-3) 3.9 3.0(-6) 6.0 2.6(-9) 8.0 1.4(-12) 10.0 3.4(-15) 9.3
128 1.3(-1) 2.0 1.2(-4) 4.0 4.7(-8) 6.0 1.0(-11) 8.0 1.5(-15) 9.9 2.2(-15) 0.7
256 3.4(-2) 2.0 7.7(-6) 4.0 7.4(-10) 6.0 3.5(-14) 8.2 1.4(-15) 0.1 4.6(-15) 1.1
Table 2: Frobenius norms and convergence rates of the final-time relative error in the numerical solution of nine and a half periods of Rabi oscillation, using method orders two through twelve.
4.1.2 Verification of Exact Gradients
We can use the Rabi oscillator in an optimal control problem by taking $\boldsymbol{\theta}=[p_{0},q_{0}]^{T}$ and $\Omega=p_{0}+iq_{0}$ so that $H_{c}=p_{0}(a+a^{\dagger})+q_{0}(a-a^{\dagger})$, and assigning some target gate $U^{\textrm{target}}\in\mathbb{C}^{2\times 2}$. We can calculate the derivatives of each entry of $U(t)$ with respect to $p_{0}$ and $q_{0}$ by applying straightforward calculus to (35):
$\frac{\partial U(t)}{\partial p_{0}}=\begin{bmatrix}-\frac{p_{0}t}{|\Omega|}\sin(|\Omega|t)&\frac{-1}{p_{0}-iq_{0}}\left(\frac{q_{0}}{|\Omega|}\sin(|\Omega|t)+ip_{0}t\cos(|\Omega|t)\right)\\ \frac{1}{p_{0}+iq_{0}}\left(\frac{q_{0}}{|\Omega|}\sin(|\Omega|t)-ip_{0}t\cos(|\Omega|t)\right)&-\frac{p_{0}t}{|\Omega|}\sin(|\Omega|t)\end{bmatrix},$ (36)
$\frac{\partial U(t)}{\partial q_{0}}=\begin{bmatrix}-\frac{q_{0}t}{|\Omega|}\sin(|\Omega|t)&\frac{1}{q_{0}+ip_{0}}\left(q_{0}t\cos(|\Omega|t)+\frac{ip_{0}}{|\Omega|}\sin(|\Omega|t)\right)\\ \frac{-1}{p_{0}+iq_{0}}\left(\frac{p_{0}}{|\Omega|}\sin(|\Omega|t)+iq_{0}t\cos(|\Omega|t)\right)&-\frac{q_{0}t}{|\Omega|}\sin(|\Omega|t)\end{bmatrix}.$ (37)
For two matrices $A,B\in\mathbb{C}^{n\times n}$, $\operatorname{Tr}(A^{\dagger}B)$ can be rewritten as
$\operatorname{Tr}(A^{\dagger}B)=\left\langle A_{\rm Re}\,,B_{\rm Re}\,\right\rangle_{F}+\left\langle A_{\rm Im}\,,B_{\rm Im}\,\right\rangle_{F}+i\left[\left\langle A_{\rm Re}\,,B_{\rm Im}\,\right\rangle_{F}-\left\langle A_{\rm Im}\,,B_{\rm Re}\,\right\rangle_{F}\right].$
The infidelity can be written as
$\mathcal{I}=1-\frac{1}{4}\left[\left(\left\langle U^{\textrm{target}}_{\rm Re}\,,U(T)_{\rm Re}\,\right\rangle_{F}+\left\langle U^{\textrm{target}}_{\rm Im}\,,U(T)_{\rm Im}\,\right\rangle_{F}\right)^{2}+\left(\left\langle U^{\textrm{target}}_{\rm Re}\,,U(T)_{\rm Im}\,\right\rangle_{F}-\left\langle U^{\textrm{target}}_{\rm Im}\,,U(T)_{\rm Re}\,\right\rangle_{F}\right)^{2}\right],$
and then the derivatives of $\mathcal{I}$ may be computed as
$\frac{d\mathcal{I}}{d{p_{0}/q_{0}}}=-\frac{1}{2}\bigg{(}\left\langle U^{\textrm{target}}_{\rm Re}\,,U_{T,{\rm Re}\,}\right\rangle_{F}+\left\langle U^{\textrm{target}}_{\rm Im}\,,U_{T,{\rm Im}\,}\right\rangle_{F}\bigg{)}\left(\left\langle U^{\textrm{target}}_{\rm Re}\,,\frac{dU_{T,{\rm Re}\,}}{d{p_{0}/q_{0}}}\right\rangle_{F}+\left\langle U^{\textrm{target}}_{\rm Im}\,,\frac{dU_{T,{\rm Im}\,}}{d{p_{0}/q_{0}}}\right\rangle_{F}\right)\\ -\frac{1}{2}\bigg{(}\left\langle U^{\textrm{target}}_{\rm Re}\,,U_{T,{\rm Im}\,}\right\rangle_{F}-\left\langle U^{\textrm{target}}_{\rm Im}\,,U_{T,{\rm Re}\,}\right\rangle_{F}\bigg{)}\left(\left\langle U^{\textrm{target}}_{\rm Re}\,,\frac{dU_{T,{\rm Im}\,}}{d{p_{0}/q_{0}}}\right\rangle_{F}-\left\langle U^{\textrm{target}}_{\rm Im}\,,\frac{dU_{T,{\rm Re}\,}}{d{p_{0}/q_{0}}}\right\rangle_{F}\right),$
where the sensitivites ${dU_{T,{\rm Re}\,/{\rm Im}\,}}/{d{p_{0}}}$ and ${dU_{T,{\rm Re}\,/{\rm Im}\,}}/{d{q_{0}}}$ may be computed explicitly using (36) and (37).
To test the correctness of our discrete adjoint implementation, we take the Hadamard gate as our target gate and again use $\Omega=p_{0}+iq_{0}=0.05(\cos(\pi/4)+i\sin(\pi/4))$ and $T=9.5\pi/|\Omega|$. We compute the gradient using the discrete adjoint method for 16 to 256 timesteps and method order 2 through 12, and compare the Frobenius norm of the difference between the analytic gradient of the continuous problem and the gradient of the discrete problem using the discrete adjoint method. The results are shown in Table 3. We observe that as the stepsize increases, the gradient converges to the analytic gradient with machine precision.
Order 2 Order 4 Order 6 Order 8 Order 10 Order 12
Steps Err Cvg Err Cvg Err Cvg Err Cvg Err Cvg Err Cvg
16 6.3 - 1.0(1) - 3.2(-1) - 4.6(-3) - 4.2(-5) - 2.6(-7) -
32 7.9 0.3 8.8(-1) 3.5 5.6(-3) 5.8 2.0(-5) 7.9 4.3(-8) 9.9 6.6(-11) 11.9
64 1.2(1) 0.6 5.8(-2) 3.9 9.0(-5) 6.0 7.8(-8) 8.0 4.3(-11) 10.0 8.3(-14) 9.6
128 3.9 1.7 3.6(-3) 4.0 1.4(-6) 6.0 3.1(-10) 8.0 4.1(-15) 13.3 4.9(-14) 0.8
256 1.0 2.0 2.3(-4) 4.0 2.2(-8) 6.0 1.1(-12) 8.1 7.1(-14) 4.1 1.3(-13) 1.4
Table 3: $\ell_{2}$ norms and convergence rates of the error in the discrete-adjoint-computed gradient of the infidelity for the Rabi oscillator problem with the Hadamard gate as the target. Numerical solution done over nine and a half periods of Rabi oscillation, using method orders two through twelve.
4.2 Speedup Experiments on a Multi-Qudit Gate Design Problem
As a practical example, we consider a two qudits coupled to a resonator bus. Using the model (4), the drift Hamiltonian is
[…]
In the model, the resonator is treated no differently than another qudit, but we use the index $R$ to indicate its significance as a resonator: the two qudits used for computation are weakly coupled to one another, but more strongly coupled to the resonator, which allows the resonator to mediate the interaction between the two qudits. For the physical parameters, we use the values from the CNOT3 example of Juqbox.jl [2], which we provide in Table 4. We also mimic Juqbox.jl’s strategy of weighting the guard penalty for the guard states exponentially based on the distance from the highest level guard state. Specifically, the weight assigned to a guard state is $0.001^{n_{G}}/N_{G}$, where $n_{G}$ is the minimum number of levels difference between the the current state and the highest guard state in each subsystem (e.g. for the tripartite system, for the state $|100\rangle$ we get $n_{G}=8$, and for $|121\rangle$ we get $n_{G}=0$), and $N_{G}$ is the total number of guard levels in the computational basis (i.e., the dimension of the guard subspace), or $N_{G}=79$ for the tripartite system. This allows optimization to populate the lower levels of the cavity, which may which advantageous for facilitating the interaction of the two qudits, while more heavily discouraging the population of the guard states which push the limits of the model using a truncated Hilbert space.
$\omega_{1}$ $\omega_{1}$ $\omega_{R}$ $\xi_{1}$ $\xi_{2}$ $\xi_{R}$ $\xi_{21}$ $\xi_{R1}$ $\xi_{R2}$
Value$/2\pi$ (GHz) 4.11 4.82 7.84 1.20(-2) 2.25(-1) 2.83(-5) 1.00(-6) 2.49(-3) 2.52(-3)
Table 4: Physical parameters used in (38) to model the Hamiltonian of two qubits coupled to a resonator bus, rounded to three significant digits. The qubits are weakly coupled to one another, but more strongly coupled to the resonator.
In the rotating frame (in which we will perform all numerical calculations), the drift Hamiltonian becomes
$H_{d}=-\frac{\xi_{1}}{2}a_{1}^{\dagger}a_{1}^{\dagger}a_{1}a_{1}-\frac{\xi_{2}}{2}a_{2}^{\dagger}a_{2}^{\dagger}a_{2}a_{2}-\frac{\xi_{R}}{2}a_{R}^{\dagger}a_{R}^{\dagger}a_{R}a_{R}-\xi_{21}a_{2}^{\dagger}a_{2}a_{1}^{\dagger}a_{1}-\xi_{R1}a_{R}^{\dagger}a_{R}a_{1}^{\dagger}a_{1}-\xi_{R2}a_{R}^{\dagger}a_{R}a_{2}^{\dagger}a_{2},$
and after making the rotating wave approximation, the control Hamiltonians are
$H_{c,K}(t;\boldsymbol{\theta})=p_{K}(t;\boldsymbol{\theta})(a_{K}+a_{K}^{\dagger})+iq_{K}(t;\boldsymbol{\theta})(a_{K}-a_{K}^{\dagger}),\quad K\in\{1,2,R\}.$
We model each of the two qubits as four-level subsystems, and the resonator as a ten-level subsystem. We wish to implement a CNOT gate on the two qudits, utilizing the resonator to mediate the interaction. We use two essential levels and two guard levels to the qudits, and one essential level and nine guard levels to the resonator (see Table 5).
Qudit $1$ Qudit $2$ Resonator
# Essential Levels 2 2 1
# Guard Levels 2 2 9
Table 5: Number of essential and guard levels for each subsystem in model of two qudits coupled to a resonator bus.
The gate is given in the laboratory frame in bra-ket notation by
$\textrm{CNOT}=|000\rangle\langle 000|+|010\rangle\langle 010|+|100\rangle\langle 110|+|110\rangle\langle 100|.$
Qudit 2 flips between the ground and excited states when qudit 1 is in the excited state. When qudit 1 is in the ground state, qudit 2 remains in the same state. The resonator always remains in the ground state (at the final time).
We fix the duration of the gate to 550 nanoseconds. As our control pulse ansatz, we use carrier waves multiplying degree 14 B-spline curves with 15 B-spline wavelets for the real and imaginary parts of the envelope, for a total of $3\times 3\times 2\times 15=270$ control parameters. For each control Hamiltonian corresponding to a subsystem, we use three carrier wave frequencies: 00, $-2\pi\xi_{1}$, and $-2\pi\xi_{2}$ for qudits $1$ and $2$, and 00, $-2\pi\xi_{R1}$, and $-2\pi\xi_{R2}$ for the resonator. The B-spline degree of 14 ensures that the Hamiltonian is smooth enough that the methods up to order 12 will exhibit their full order of accuracy in the asymptotic region. For the lower-order methods, we could lower the degree of the B-spline curves to reduce the cost of evaluating the control functions while keeping the full order of accuracy, but we use the same degree for all orders for the sake of keeping the comparison as direct as possible.
4.2.1 Determining Appropriate Step Size
In order for the optimization to find control pulses which implement a gate with low infidelity, the numerical solution of Schrödinger’s equation must be accurate enough to ensure that the error in the infidelity is reasonably smaller than the infidelity itself. With this in mind, before performing the optimization using the Hermite method of order $N$ we perform a numerical experiment to determine an appropriate stepsize $\Delta t$ to use in order to reach a target relative error in the numerical solution of Schrödinger’s at the final time of the gate.
For each order of the method, starting with two timesteps and doubling the number of timesteps each iteration until the state history cannot fit into main memory, we perform a forward evolution, record the final state, perform the rest of the gradient computation (adjoint evolution and gradient accumulation), and record the time taken for the full gradient computation. We do this for 25 samples of the control vector $\boldsymbol{\theta}$ from a uniform distribution with a range of $[-0.05,0.05]$ for each control parameter, which enforces a maximum amplitude of 50 MHz on the B-spline envelope for each carrier wave for each control pulse. For each order of the method, the same 25 samples are used, so that each method is attempting to simulate the exact same analytic dynamics. An absolute tolerance of $10^{-15}$ was for the linear solves, which were performed using GMRES. This tolerance could be relaxed for greater computational efficiency, depending on the accuracy required of the state vector. In Figure 3, we report the number of timesteps, mean time to compute the gradient, and mean relative error at the final time for each order of the method. The samples were collected using a single core of a Intel® Xeon® Gold 6148 CPU clocked at 2.40GHz, and 32 GB of memory, in Michigan State University’s high-performance computing cluster.
Figure 3: Comparison of the number of timesteps used, the time taken to compute the gradient, and the relative error in the numerical approximation of the state, for the three subsystem example, starting in the initial states of the gate design problem: $|000\rangle,|001\rangle,|010\rangle$, and $|011\rangle$. The HOHO method is used with orders 2, 4, 6, 8, 10, and 12. We also compare with the second-order method Störmer-Verlet, as implemented in the Julia package Juqbox.jl.
In order to estimate the number of timesteps needed to reach a target relative error, we use linear interpolation of the number of timesteps and the mean relative error in the case that we have recorded mean relative errors above and below the target relative error. Otherwise, we estimate the number of timesteps using a least squares fit of the the number of timesteps and the mean relative error in the region of asymptotic convergence.
This heuristic approach warrants some skepticism that we can use a single stepsize in order to consistently achieve a relative error below the target relative error during the optimization, since the control vector (and hence the dynamics) changes with each iteration of the optimization. However, in our experiment we observe that the standard deviation of the relative error is always an order of magnitude smaller than the mean relative error, indicating that the magnitude of the relative error is consistent for different values of $\boldsymbol{\theta}$. We use this numerical observation to justify using a single stepsize during the optimization. Table 6 shows the number of timesteps required for to reach several target relative final-time errors for each order of the Hermite method, and Table 7 shows the expected speedup of the gradient computation (to perform forward evolution, adjoint evolution, and gradient accumulation) for each order of the HOHO method compared to Störmer-Verlet, as implemented by Juqbox.jl, based on the average time taken to compute the gradient over the 25 different control vectors.
Number of Timesteps Required
Target State Error Order 2 (SV) Order 2 Order 4 Order 6 Order 8 Order 10 Order 12
1.0(-1) 5,287 7,401 822 378 182 175 92
1.0(-3) 57,051 75,775 3,530 1,151 609 388 286
1.0(-5) 540,005 778,948 11,198 2,505 1,104 642 432
1.0(-7) 5,334,675 8,028,679 35,328 5,409 1,974 1,029 644
Table 6: The number of timesteps for each order of the Hermite method (and for Störmer-Verlet) required to reach a target relative final-time error in the numerical solution of Schrödinger’s equation for the three subsystem example, starting in the initial states of the gate design problem: $|000\rangle,|001\rangle,|010\rangle$, and $|011\rangle$. Results are based on the average accuracy for 25 randomly sampled control vectors.
Speedup Factor
Target State Error Order 2 (SV) Order 2 Order 4 Order 6 Order 8 Order 10 Order 12
1.0(-1) 1.0 0.4 0.7 0.2 0.1 0.1 0.1
1.0(-3) 1.0 0.5 3.0 0.9 0.5 0.3 0.2
1.0(-5) 1.0 0.5 11.0 12.8 2.7 1.9 1.1
1.0(-7) 1.0 0.5 36.9 87.3 54.2 16.8 11.7
Table 7: The speedup factor for each order of the Hermite method over Störmer-Verlet for the three subsystem example, starting in the initial states of the gate design problem: $|000\rangle,|001\rangle,|010\rangle$, and $|011\rangle$. Results are based on the average time elapsed for 25 randomly sampled control vectors.
In Tables 8 and 9, we again show the number of timesteps required and the average speedup for the three subsystem example, but for the evolution of a single initial state: $|933\rangle$, the most highly excited state of the system, for which the dynamics are expected to happen over much shorter timescales. This represents a worst-case scenario computationally, as more timesteps are required to resolve the faster dynamics, and consequently the higher-order methods are much more efficient, since their accuracy more significantly with the number of timesteps. Fast dynamics could also occur even in the computational subspace when there are many qudits, for example, or if the computation is done in the laboratory frame instead of the rotating frame.
More analytic approaches for determining the number of timesteps can be used, such as an analysis of the smallest frequency of the dynamics based on the eignevalues of the Hamiltonian [2], but this numerical experiment is simple and captures the effect of changing the control vector on the numerical accuracy, resulting in a robust estimate of the number of timesteps needed to reach a particular level of accuracy. And although the experiment requires performing dozens of gradient computations, it is still cheap compared to the control optimization, which may perform hundreds or thousands of gradient computations.
Number of Timesteps Required
Target State Error Order 2 (SV) Order 2 Order 4 Order 6 Order 8 Order 10 Order 12
1.0(-1) 232,401 311,607 14,684 4,495 2,292 1,424 1,041
1.0(-3) 2,004,693 2,267,052 46,606 9,755 4,136 2,351 1,556
1.0(-5) 17,452,973 16,493,625 146,904 21,050 7,381 3,754 2,316
1.0(-7) 151,946,628 119,997,072 463,615 45,367 13,143 5,967 3,413
Table 8: The number of timesteps for each order of the Hermite method (and for Störmer-Verlet) required to reach a target relative final-time error in the numerical solution of Schrödinger’s equation for the three subsystem example, starting in the most highly excited state: $|933\rangle$. Results are based on the average accuracy for 25 randomly sampled control vectors.
Speedup Factor
Target State Error Order 2 (SV) Order 2 Order 4 Order 6 Order 8 Order 10 Order 12
1.0(-1) 1.0 0.7 5.0 5.5 1.2 0.9 0.8
1.0(-3) 1.0 0.9 15.4 26.9 25.1 9.3 6.3
1.0(-5) 1.0 1.2 45.1 121.4 142.9 101.4 52.5
1.0(-7) 1.0 1.5 136.7 530.1 774.6 702.4 459.1
Table 9: The speedup factor for each order of the Hermite method over Störmer-Verlet for the three subsystem example, starting in the most highly excited state: $|933\rangle$. Results are based on the average time elapsed for 25 randomly sampled control vectors.
4.2.2 Performing Gate Optimization
We now perform the gate optimization. We compare the performance of each order of the Hermite method for target relative final-time errors of $10^{-1}$, $10^{-3}$, $10^{-5}$, and $10^{-7}$. The optimization is managed by IPOPT (Interior Point OPTimizer) [39], a software library for solving large scale nonlinear optimization problems using a primal-dual interior point method. We configure IPOPT to perform a limited-memory quasi-Newton approximation of the Hessian of the objective function for use in line searches (second-order optimization). IPOPT also allows for box constraints on the control vector, which we use to enforce our constraint on the control pulse amplitudes.
For each order and target error, we perform the optimization from ten initial control vectors, sampled from a uniform random distribution with a range of $[-0.005,0.005]$. The upper and lower bounds on each control parameter (enforced by IPOPT) are set to ±$\pm 0.05$, enforcing a maximum amplitude of 50 MHz on the B-spline envelope of each carrier wave for each control pulse, the same as the upper and lower limits on the random distribution in the stepsize experiment. The same ten initial control vectors are used for all orders and all target errors, so that every optimization starts from the same analytic control pulses. Each optimization is allowed to run for six hours before being halted.
In Figure 4, for each target error we plot the generalized gate infidelity against the number of IPOPT iterations completed for each order of the Hermite method. In the figure, we do not plot the “watchdog” iterations, in which the optimizer aggressively explores new search directions, resulting in large changes in the objective function which would make the graph difficult to interpret (but the watchdog iterations are still counted in the total). These samples were collected using the same setup as the stepsize test, except we use four CPU cores instead of one and use multithreading to parallelize the forward and adjoint evolutions over the different initial and terminal conditions. In Figure 5, we plot the time evolution of the population of the computational basis states for the control found that best minimized the generalized gate infidelity for a target final state error of $10^{-7}$.
Figure 4: Optimization of a CNOT gate using two qudits coupled to a resonator bus as the mode, using B-spline curves multiplied carrier waves as the control pulse ansatz, using the methods and numbers of timesteps prescribed by Table 6 in order to reach target relative final state errors on the orders of $10^{-1}$, $10^{-3}$, $10^{-5}$, and $10^{-7}$.
Figure 5: Time evolution of the state populations for the CNOT gate by the best (lowest generalized infidelity) control pulses found by the optimizations performed with a target relative final state error of $10^{-7}$ for the two qudits plus resonator model (see Figure 4).
4.2.3 Analysis
From the the plot of the number of the mean relative error against the number of timesteps for the three subsystem example in Figure 3, we see that each order of the Hermite exhibits the full order of convergence. From the plot of the mean gradient computation time against the number of timesteps, we can see that the computation time increases with the number of timesteps, and the asymptotic behavior of each order of the method is the same, indicating that for the computation time of the higher-order methods is roughly a constant multiple of the computation time of the lower-orders, which is to be expected. Interestingly, for the sixth-twelfth-order methods the computation time decreases around $2^{9}-2^{11}$ timesteps, for which the stepsize ranges from $1.07\geq\Delta t\geq 0.27$ nanoseconds. We attribute this to a decrease in the number of GMRES iterations required in the linear solves of the forward and adjoint evolutions due to the effectiveness of our preconditioner and our initial guess (using Taylor expansion) when $\Delta t<1$ nanosecond.
Most importantly, the plot of the mean gradient computation time against the mean relative error can be used to estimate which method order will be the fastest for a particular target error. For example, for a target error of $10^{-3}$, the gradient computation is fastest for the fourth-order method, but for a target error of $10^{-6}$ the sixth-order method is fastest. This also demonstrates the effectiveness of high-order methods; for a target error less than $10^{-2}$, the methods of order four and higher are always faster than the second-order methods.
The computational efficiency of the high-order methods is showcased by Tables 7 and 9. When using the initial conditions from the gate design problem and trying to reach final state errors of $10^{-3}$-$10^{-5}$, using high-order methods results in average speedup factors of approximately $3$-$13$ compared to Störmer-Verlet (when choosing the most efficient order of the method for each target error). When the initial condition is the highly excited state $|933\rangle$, the average speedup factors are in the range $27$-$143$!
In addition to prescribing the number of timesteps required to reach a target error, Tables 6 and 8 also highlight another advantage of the high-order methods: reduced memory requirements. The memory required to store the state history in the forward and adjoint evolutions scales linearly in the number of timesteps and the dimension of Hilbert space. Because the high-order methods can obtain highly accurate solutions with a small number of timesteps, we are often able to store the entire state history in memory. This allows us to avoid using the “memory-lean” version of the gradient calculation, which would require more computation and produce less accurate gradients.
Figure 4 highlights the effectiveness of high-order methods for the optimization process. We observe that for different order methods and different target errors, but starting from the same initial control vector, the generalized gate infidelity evolves roughly the same as a function of the number of IPOPT iterations. So our ability to optimize the objective function is determined mainly by the number of IPOPT iterations we are able to perform. Consequently, a method that can compute the gradient faster and therefore perform more IPOPT iterations in a fixed amount of wall time can converge to a lower objective function value. Because the higher-order methods can typically compute the gradient faster, they are able to optimize the objective function to a lower value in the same amount of time. As the target error decreases, the the higher-order methods become more effective.
5 Conclusions and Future Work
In this work, we derived a discrete adjoint method for efficiently computing gradients of objective functionals depending on the number solution of a linear system of ODEs solved by Hermite Runge-Kutta methods. Furthermore, we presented techniques for improving the computational efficiency of the method when the time-dependence of the Hamiltonian takes the typical form for systems controlled by control pulses, along with an efficient preconditioner for models where the drift Hamiltonian is diagonal.
We tested our method on the problem of designing a CNOT gate for two qudits coupled to a resonator bus, with guard levels included to ensure the validity of the model. We observed significant improvement in time and memory usage when using Hermite Runge-Kutta methods with the discrete adjoint method derived in this work (as implemented by QuantumGateDesign.jl) and comparing to the Störmer-Verlet method (as implemented by Juqbox.jl). When computing gradients to moderate levels of accuracy (relative final state error $10^{-3}$-$10^{-5}$) for the “worst-case” scenario that the system starts in its most highly excited state, we observed speedup factors of 27-143 and memory reduction factors of approximately $1,200$-$7,500$ by using high-order Hermite Runge-Kutta methods.
For a more strict relative final state error requirement of approximately $10^{-7}$, a speedup factor of 775 and a memory reduction factor of 11,561 was achieved by the most time-efficient method, and a speedup factor of 460 and a memory reduction factor of 44,520 was achieved by the most memory-efficient method. For stiffer systems, we expect high-order methods to be even more effective.
Future work will focus on both improving the efficiency of our method, and on using the efficiency and flexibility of our method to explore new avenues in quantum optimal control. To improve the efficiency of our method, we could develop better preconditioners for solving the linear system of equations (16) at each timestep, especially for models with a non-diagonal drift Hamiltonian. We could also explore adjusting the number of timesteps and the method order to increase the accuracy as the optimization progresses, maximizing efficiency. A new application of the method could be to optimize control pulses which include the response function of the hardware used to generate them in the model.
6 Acknowledgments
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No 2235783 (SL) and by NSF under Grant Number 2436319 (DA). Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Part of this research was performed while one of the authors (SL) was visiting the Institute for Pure and Applied Mathematics (IPAM), which is supported by the National Science Foundation (Grant No. DMS-1925919). This material is based upon work supported by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research (ASCR), under Award Number DE-SC0025424 (DA). This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof. We acknowledge Michigan State University’s support through computational resources provided by the Institute for Cyber-Enabled Research. Thanks to the creators of the Julia programming language [4] and of the packages used in the creation of this paper, including Makie.jl [10], DataFrames.jl [5], and others which do not request citations in academic literature.